A refrigerator has to transfer an average of 263 J of heat per second from temperature -10 C to 25 C. Calculate the average power consumed assuming ideal reversible cycle and no other losses.

Grade:12

1 Answers

Pramod J AskiitiansExpert-IIT-B
13 years ago

Dear student,

\ COP = \frac{T_L}{T_H - T_L}

 

=> COP = 263/35 = 7.514 = QH/W

=> W = 263/7.514 = 35 Watts